# Generative Large Language Models in Automated Fact-Checking: A Survey

### v1

@@ Appendix C Model Analysis
    The GPT Family. Generative Pre-trained Transformers (GPT) is a family of LLMs developed by OpenAI and includes mostly closed-source models, accessible by paid API. This category includes all GPT models (e.g. GPT-3, GPT-3.5, or GPT-4), CODEX. In addition to these closed-source models, there are also several open-sourced models, such as the GPT-Neo model.
@@
Figure 3 depicts LLMs employed in more than five papers, accompanied by the respective count of papers each model utilized. A total of 35 various LLMs were employed in the collection of 70 papers.
@@
Name     Check-worthy claim detection     Previously fact-checked claims detection     Evidence retrieval     Fact verification # Lang. Size Citation
@@
COVID-19 Scientific 1 142 Lee et al. (2020)
COVID-Fact 1 4k Saakyan et al. (2021)
ChatGPT-FC 1 22k Li et al. (2023b)
Check-COVID 1 2k Wang et al. (2023a)
ClaimBuster 1 24k Arslan et al. (2020)
ClaimDecomp 1 1k Chen et al. (2022)
@@
Paper Approaches Techniques # Lang.
@@
Chen et al. (2022) - claim decomposition - fine-tuning 1
@@
Kamoi et al. (2023b) - text generation - few-shot 1
Li et al. (2023a) - text generation - few-shot 1
@@
Wang et al. (2023b) - 3-class classification - zero-shot 1
Sundriyal et al. (2023) - text generation - few-shot - zero-shot CoT 1
@@
Table 2: Work addressing the task of check-worthy claim detection in the context of fact-checking. Each paper includes the utilized approaches, techniques and the number of languages analyzed.
Paper Approaches Techniques # Lang.
@@
Paper Approaches Techniques Knowledge base # Lang.
@@
Wang et al. (2023a) - rationale selection - few-shot - scientific papers 1
@@
to lllll Paper Approaches Techniques Used Evidence # Lang.
@@
Wang et al. (2023a) - 3-class classification - few-shot - open-book 1
@@
Hu et al. (2023) - 3-class classification - zero-shot - few-shot - zero-shot CoT - few-shot CoT - self-refinement - self-consistency - open-book 27
@@
Russo et al. (2023) - explanation generation - fine-tuning - open-book 1

### v2

@@ Appendix D Model Analysis
    The GPT Family. Generative Pre-trained Transformers (GPT) is a family of LLMs developed by OpenAI and includes mostly closed-source models, accessible by paid API. This category includes all GPT models (e.g. GPT-3, GPT-3.5, or GPT-4) or CODEX. In addition to these closed-source models, there are also several open-sourced models, such as the GPT-Neo model.
@@
Figure 2 depicts LLMs employed in more than five papers, accompanied by the respective count of papers each model utilized. A total of 33 various LLMs were employed in the collection of 69 papers.
@@
Name     Claim detection     Previously fact-checked claims detection     Evidence retrieval     Fact verification # Lang. Size Citation
@@
Paper Methods Techniques # Lang.
@@
Chen et al. (2022) - question generation - fine-tuning 1
@@
Kamoi et al. (2023b) - claim generation - few-shot 1
Li et al. (2023a) - claim generation - few-shot 1
@@
Sundriyal et al. (2023) - claim generation - few-shot - zero-shot CoT 1
@@
Table 3: Work addressing the task of claim detection in the context of fact-checking. Each paper includes the utilized approaches, techniques and the number of languages analyzed.
Paper Methods Techniques # Lang.
@@
Paper Methods Techniques Knowledge base # Lang.
@@
Wang et al. (2023b) - rationale selection - few-shot - scientific papers 1
@@
to lllll Paper Methods Techniques Used Evidence # Lang.
@@
Wang et al. (2023b) - 3-class classification - few-shot - open-book 1
@@
Russo et al. (2023a) - data generation - zero-shot 1
Russo et al. (2023b) - explanation generation - fine-tuning - open-book 1